OpenStudy (anonymous):

eval the integral integrate x^3 (sqrt x^4+2) dx

OpenStudy (anonymous):

I think a U-substitution will work here.

OpenStudy (anonymous):

Is the +2 also in the square root?

OpenStudy (anonymous):

Try substituting u=x^4+2

OpenStudy (anonymous):

yes 2 is also in sort

OpenStudy (anonymous):

can you give me step by step so I know exactly what to di cuz I'm a little confused w/ U sub

OpenStudy (anonymous):

Ah, ok. The first step is to see that you have both x^3 and x^4 in the function and 4x^3 is the derivative of x^4. If you let u=x^4+2, then du = 4x^3 dx The function you are integrating is then \(\large (1/4)u^{1/2} du \)

OpenStudy (anonymous):

how did you get 1/4?